OpenStudy (ashlyhall8):

Can you have a right triangle where the lengths of the legs are whole numbers and the length of the hypotenuse is 23? Explain.

OpenStudy (welshfella):

in a right angled triangle , by Pythagoras. c^2 = a^2 + b^2 where c is the hypotenuse and a and b are the 2 legs c ^2 = 23^2 = a^2 + b^2 529 = a^2 + b^2

OpenStudy (welshfella):

now you need to find 2 perfect square numbers which will add up to 529. Do you think there exist?

OpenStudy (ashlyhall8):

Idk i just need the answer so i can explain

OpenStudy (mjdennis):

He's trying to help _you_ find the answer so you _can_ explain!

OpenStudy (ashlyhall8):

But idk how

OpenStudy (welshfella):

If you catn find any then we know that there are no whole numbers which fit the triangle.

OpenStudy (ashlyhall8):

im doing my classes online and i have no clue what im doing

OpenStudy (welshfella):

to be honest I cant think of a quick way to do this other than to subtract square numbers from 529 and test if the result is also a square number for example assume that one leg is 2 so the square is 4 529 - 4 = 525 Now 525 is not a square number (tested with calculator)

OpenStudy (ashlyhall8):

so how do i do it

OpenStudy (welshfella):

next try 3^2 = 9 529 - 9 = 520 - nit a square 4^2 = 15 529 - 16 = 513 not a square

OpenStudy (welshfella):

- a long process Maybe theres a quicker way...

OpenStudy (ashlyhall8):

ok. im just gonna guess

OpenStudy (welshfella):

I'm pretty sure you ca'nt have a triangle with legs = whole numbers if hypotenuse = 23.

OpenStudy (mjdennis):

A spreadsheet says @welshfella is correct, but the only explanation I know involves checking all the combinations. Good luck.

OpenStudy (welshfella):

yea I cant think of another way either.

OpenStudy (welshfella):

you'd have to check all square numbers up the the one after 529/2